TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Hugh S. Gregory
date: 2003-03-19 23:25:00
subject: 3\08 Pt-1 ISS On Orbit Status 08-03-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
3\08 ISS On-Orbit Status 08-03-2003
Part 1 of 2

ISS On-Orbit Status 3/8/03

All ISS systems continue to function nominally, except as noted 
previously or below. Another Saturday for "Dr. Scary Science" (Awed 
Flight Controllers' title for Science Officer Dr. Pettit and his 
amazing "Saturday Morning Science" show downlink).

The ISS residents completed the regular weekly 3-hr. house cleaning,
designed to keep their heavenly abode spic and span. ["Uborka 
stantsii" features removal of food waste products, cleaning of 
compartments with vacuum cleaner, wet cleaning of surfaces with 
disinfectants and cleaning of fan screens to avoid temperature rises.]

FE-1 Nikolai Budarin conducted his daily checkup and watering of the 
Russian BIO-5 Rasteniya-2/Lada-2 ("Plants-2") plant growth experiment.

For FE-2 Don Pettit, MSG (Microgravity Science Glovebox) went into 
another day, progressing at a carefully charted snail pace. [Initial 
indications of the latest power re-connect (P121) showed no 
off-nominal current draw on ESEM3 (exchangeable standard electronic 
module 3), exonerating its power path as a possible source of the 
original current-trip anomaly. IF ESEM3 remains nominal with the 
currently installed P106 connector (for Work Volume temperature, 
humidity, and Delta-P sensors), P101 will be reconnected next, 
powering the PCV (process control valve) motor limit switch. 
Troubleshooting will continue into next week. At some point while the 
rack is powered up and Ku-band is on, Pettit will conduct a simple 
"shake test", with actions like fully extending the Core Facility and 
opening and closing the video drawer.]

Nick took care of the daily routine maintenance of the SOSh life 
support system, incl. ASU toilet subsystem, and Don prepared the IMS 
inventory database for automated file import/export.

The crew also performed their daily physical exercise program on TVIS
(treadmill with vibration isolation and stabilization), RED (resistive
exercise device), CEVIS (cycle ergometer with vibration isolation), 
and VELO ergometer with load trainer.

Because of the current CEVIS control panel failure, the crew was asked 
to perform next week's (3/12) periodic fitness evaluation (PFE) on the 
Russian VELO. [PFE, usually performed on the CEVIS, checks up on blood 
pressure (BP) and electrocardiogram (ECG) during programmed exercise, 
with the BP/ECG and HRM (heart rate watch) equipment.]

Because of CDR Bowersox's concern regarding potential asymmetric 
loading when reinstalling the used Flexpack canister #1010 in the RED, 
the decision was made to withdraw the idea for now, pending further 
study. [The crew can continue to use RED with the current #1001 and 
#1003 canisters.]

The VOA (volatile organics analyzer) suddenly shut down yesterday, and 
the crew was instrumental in recovering it. It is currently powered 
and in (nominal) standby mode. [Engineers are not sure what caused the 
shutdown and which troubleshooting action recovered it. Data files 
will be dumped and analyzed to determine the exact cause of the 
anomaly.]

As reported yesterday, an RGA (rate gyro assembly) "Caution" message 
popped up during the ISS attitude transition. It is now understood and 
in the process of being corrected. [The message, "RGA Dilemma Lab", 
which is still displayed by the C&W system ("latched"), was caused by 
the difference between rates measured by the two RGAs exceeding a 
pre-set threshold value in the new GNC (guidance, navigation & 
control) R3 software. The threshold, which is not "tuned" for all 
on-orbit activities, was simply too tight for the attitude change 
maneuver, and the RGAs performed nominally. Next week, a new PPL 
(pre-positioned load) software file with larger thresholds will be 
uploaded and the C&W then "unlatched". In the future, the PPL 
thresholds will be tuned from on-orbit data and then loaded.]

In response to the crew's report of a noisy IMV (intra-modular 
ventilation) fan (ref. yesterday's status), troubleshooting of Lab and 
Node air ducts, etc. was added to the "job jar" task list for 
tomorrow, and appropriate procedures were uplinked overnight. [If the 
crew chooses to perform the activity, the ground wants to be notified 
two hours in advance so the proper Flight Controllers can be in place 
to support it.]

In the continuing interest of conserving precious ISS resources, a 
procedure was uplinked for double-sided printing of any PDF or Word 
document on the onboard Epson 800 printer.

RSC-Energia and TsUP have confirmed that the portside SM Kurs 
antenna/cable survey with the SSRMS (space station remote manipulator 
system) will be conducted on 3/10 (Monday) in sunlight. [For this 
purpose, the solar array will be feathered at about 5:40pm EST, the 
start of the preceding eclipse (night period).]

ISS flight attitude continues in LVLH TEA (local vertical/local 
horizontal torque equilibrium attitude), through Progress thruster 
test/reboost on 3/12.

Expedition 6 may not include a woman, but on the ground and in space 
women are an important and highly respected part of the ISS family. To 
them: Happy "International Women's Day"!

(continued)

---
* Origin: SpaceBase[tm] Vancouver Canada [3 Lines] 604-473-9357 (1:153/719)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.